OpenStudy (anonymous):

If f(x) = 2x^2 - 30, find f(4). What do I do? Please show step by step thank you!

OpenStudy (lochana):

put 4 instead x f(4) = 2(2)^2 - 30 = 2(2)(2) - 30 = 8 - 30 = -22

OpenStudy (anonymous):

f(x) = 2x^2 - 30 f(4) = 2(4^2) - 30 f(4) = 2(16) - 30 f(4) = 32 - 30 f(4) = 2 basically, your just subbing in 4 for x

OpenStudy (anonymous):

better?

OpenStudy (anonymous):

oh I see thank you lochana! :)

OpenStudy (anonymous):

ya thank you both.

OpenStudy (anonymous):

@Animal<3er you should interact with the person helping you here.. :)

OpenStudy (zbrandz23):

:D

OpenStudy (anonymous):

that was pretty easy just forgot how to do this but, thanks to you to I remember now.